A ball is thrown vertically upwards with a velocity of 49 m/s. Calculate
the total time it takes to return to the surface of the earth.

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Time take t = ?
v = u + gt
∴ 0 = 49 + (–9.8) × t
∴t = -49/-9.8
= 5 s
Total time taken to return the surface of the earth by the ball is 5 s + 5 s = 10 s.

Thanks